OpenStudy (anonymous):

Write the equation of a line given the points (2, 5) and (-1, -2). HOW do you get this answer from that? y=7/3x+1/3 Please do explain :) Thanks :)

OpenStudy (helder_edwin):

first find the slope between the two points.

OpenStudy (anonymous):

I did that but for some reason it came out completely wrong haha Linear equation completely confuse me :/

OpenStudy (anonymous):

Dont I use y2-y1 M= --------- ???? x2-x1

OpenStudy (helder_edwin):

the slope is between (2,5) and (-1,-2) \[ \large m=\frac{5-(-2)}{2-(-1)}=\frac{7}{3} \]

OpenStudy (helder_edwin):

did u get the same?

OpenStudy (anonymous):

OH! I see what you did. Yes I now got that

OpenStudy (helder_edwin):

now given the slope \(m\) and a point \((x_0,y_0)\) (which is one of the points u r given) the equation of the line is \[ \large y-y_0=m(x-x_0) \] just plug-in the values.

OpenStudy (anonymous):

so this is a point slope equation? yes?

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

Thats probably where I got screwed up :) Thank you soooo much! I might ask your help when I get stuck =P haha

OpenStudy (helder_edwin):

u r welcome

OpenStudy (anonymous):

I became a fan so Ill send you a message if I need some help? If you dont mind? :) You actually explain compared to most on this site >.<

OpenStudy (helder_edwin):

sure thing.

OpenStudy (anonymous):

Great :)

OpenStudy (helder_edwin):

right now i have to go (lunch time here). good luck.

OpenStudy (anonymous):

Thanks again! :)